# Professionally translated parallel data for 115 under-represented languages

### v1

@@ Abstract
We open-source Smol (Set of Maximal Overall Leverage)111huggingface.co/datasets/google/smol, a suite of training data to unlock translation for low-resource languages (LRLs). Smol has been translated into 115 under-resourced languages, including many for which there exist no previous public resources, for a total of 6.1M translated tokens. Smol comprises two sub-datasets, each carefully chosen for maximum impact given its size: SmolSent, a set of sentences chosen for broad unique token coverage, and SmolDoc, a document-level source focusing on a broad topic coverage. They join the already released Gatitos for a trifecta of paragraph, sentence, and token-level content. We demonstrate that using Smol to prompt or fine-tune Large Language Models yields robust ChrF improvements. In addition to translation, we provide factuality ratings and rationales for all documents in SmolDoc, yielding the first factuality datasets for most of these languages.
@@
Isaac Caswell∗ Elizabeth Nielsen∗ Jiaming Luo Colin Cherry Geza Kovacs Hadar Shemtov Partha Talukdar Dinesh Tewari Baba Mamadi Diane Koulako Moussa Doumbouya Djibrila Diane Solo Farabado Cissé Google {Research, Deepmind} Stanford University NKo USA INC ∗{icaswell, eknielsen}@google.com
@@ 1 Introduction
Most of the world’s 7000 or so languages have few or no resources available for machine translation. However, for an expensive task like professional translation, it is not clear how to make the best use of a limited budget to generate the most effective data. As shown by the Gatitos dataset (Jones et al., 2023), training data consisting of individually translated tokens without context provides large benefits to translation quality for low-resource languages at the lowest cost. However, gains quickly saturate, as single tokens are not very expressive. Sentence-level data is better for a model once token-level data saturates, but it has much more inherent redundancy; and document-level data is even more effective…and more redundant. A resource like Gatitos is a useful first step for the languages it covers, but without sentence- or document-level resources, there is a limit to how effective translation models can be.
In this work, we release the Smol dataset, which provides professionally translated sentence- and document-level data for 115 LRLs, supplementing the token-level translations from Gatitos. Smol contains two sub-datasets:
SmolSent: 863 English sentences covering 5.5k of the most common English tokens,222In this paper, ‘token’ refers to logical/linguistic words, not subword tokens from a model’s vocabulary. professionally translated into 81 languages.
SmolDoc 584 English documents covering a wide range of topics, domains, and tokens, generated by a LLM and professionally translated into 100 languages.
We demonstrate the utility of these data for finetuning and prompting LLMs for translation, and also provide factuality annotations for all documents.
@@ 2 Related work
There are several training datasets that comprise parallel data for Low-Resource Languages (LRLs), which we operationally define as any language beyond the first 100 or so supported by most traditional crawls and MT providers. The NLLB-Seed dataset is a sentence-level training set consisting of 6k sentences selected from English Wikipedia and professionally translated into 39 LRLs Team et al. (2022). The Gatitos dataset Jones et al. (2023) represents another approach, consisting of a 4000-entry lexicon, translated into 170 LRLs. There are also several professionally-translated evaluation sets, namely Flores-101 and Flores-200 Goyal et al. (2022); Team et al. (2022), and NTreX (Federmann et al., 2022a).
@@
While highly multilingual, professionally translated training data is rare, there is a growing number of bottom-up community datasources organized through research collectives like Masakhane ( et al., 2020), Turkish Interlingua (Mirzakhalov et al., 2021a, b), and GhanaNLP (Azunre et al., 2021a); and conferences and workshops like AfricaNLP, AmericasNLP (Mager et al., 2021) and ArabNLP. These datasets are usually generated by researchers fluent in the languages, and are therefore especially high quality. In addition to providing datasets, such efforts frequently also provide models and baselines, or even public interfaces, like the Khaya Translator Web App333https://ghananlp.org/project/translator-webapp/ by GhanaNLP for West African languages, and the lesan.ai444https://lesan.ai/translate translation website for Ethiopian languages.
@@
Participation is especially strong from the African continent, including corpora and models for pan East-African languages (Babirye et al., 2022), languages from the Horn of Africa (Hadgu et al., 2022), Ethiopian languages (Teferra Abate et al., 2018; Gezmu et al., 2021), Ugandan languages (Akera et al., 2022), Emakhuwa (Ali et al., 2021), South-African languages (Eiselen and Puttkammer, 2014), Setswana and Sepedi (Marivate et al., 2020), Yorùbá (Adelani et al., 2021b, a), Oshiwambo (Nekoto et al., 2022), Igbo (Ezeani et al., 2020), Zulu (Mabuya et al., 2021), Twi (Azunre et al., 2021b), Gbe (Hacheme, 2021), Bambara (Tapo et al., 2021), and Fon (Emezue and Dossou, 2020). Outside of Africa, corpora have been created for languages of the Americas, including for four indigenous languages of Peru in Bustamante et al. (2020), the numerous results on the largely South- and Central American languages from the first AmericasNLP conference (Mager et al., 2021), and the Inuktitut language of Canada (Joanis et al., 2020). Datasets for lower-resourced languages of India have also sprung up, including the 13-language PMIndia (Haddow and Kirefu, 2020), and datasets focused on languages of the Northeast like Mizo (Thihlum et al., 2020), Khasi (Laskar et al., 2021) and Assamese (Laskar et al., 2020). Finally, a variety of such datasets and models are available for public use on HuggingFace555https://huggingface.co/datasets?multilinguality=multilinguality:translation&task_categories=task_categories:translation&sort=downloads or Zenodo666https://zenodo.org/communities/africanlp/.
3 Methods for getting text to translate
Translation requires significant investment, so great care needs be put into choosing good/effective sentences to translate. This problem is especially tricky because one can’t iterate on the method: once text is translated, the budget is exhausted and one can’t simply select a new source text. It is possible to run ablations from high resource languages, as described in the next section, but it is questionable how much such approaches can actually generalize to true Low-Resource Languages.
For both SmolDoc and SmolSent (as with Gatitos), selection/generation of source text is done in English. This has obvious biases, including the danger of translationese output, a scarsity of translators fluent in English, and a lack of culturally specific domains, but also has large advantages—most notably better scaling for quality control, and the ability to do Researcher in the Loop.
@@ 3.1 SmolSent: Token Set Cover
Our basic motivation for creating SmolSent was to help models overcome vocabulary issues, which are common for the lowest-resource languages (Bapna et al., 2022). the alligator problem Nielsen et al. (2025). Therefore, we frame this as a set-cover problem, and pick the smallest set of sentences that cover the largest set of target tokens. The tokens we chose to cover (the target set) were the English side of Gatitos, as well as the most common 2,500 tokens from an English web crawl. Set cover is NP-hard, so we approximate it with a greedy algorithm that iteratively picks the sentence with the highest coverage percent, defined as the percentage of its tokens that are in the target set.
3.1.1 Preliminary work on Token Set-Cover
@@
To evaluate the token set-cover approach, we started by selecting data from existing web-scraped parallel data. We pretrain a multilingual NMT model on parallel data from 294 language pairs from Madlad-400 Kudugunta et al. (2023), with nine languages held out to simulate LRLs. We fine-tune this model on sets of existing parallel data in each of the held-out languages, and evaluate on Flores-200. Details on the experimental set-up can be seen in Appendix A.1.
In addition to Greedy Token Set-Cover, we explore two methods that balance data diversity and data quality. First, we implement Ambati et al. (2011)’s ‘density-weighted diversity’ (DWD) metric, which is an n-gram based metric for diversity and quality. Second, we implement an embedding-based version of DWD, which takes the weighted harmonic mean of perplexity under the Palm 2 model Anil et al. (2023) (proxy for quality), and embedding distance on mBERT sentence embeddings (proxy for diversity). We apply both methods to the English side of the parallel data only, to simulate the case where we don’t yet have LRL translations. As a baseline, we randomly select sentences.
@@
3.1.2 Researcher in the Loop (RITL)
Despite its success in the ablation, Greedy Token Set Cover had several problems when we scaled it to select from among all the English sentences of CommonCrawl. Firstly, it is maximized by honeypots, or nonsense strings dense in content words (Appendix Table 4); and secondly, it biases towards short sentences, causing length distribution artifacts
These are tricky problems to solve in the general case. However, a dataset like Smol is small enough to manually inspect. Therefore we develop Researcher in the Loop Greedy Set-Cover (Algorithm A.2), where the domain expert (the researcher) can look at and edit each individual sentence.777This work was conducted before the advent of LLMs. Today, this could be simplified using LLMs as autoraters. The result of this process is SmolSent, a set which uses 863 sentences to cover 5519 unique tokens. Qualitatively, SmolSent consists of complex sentences with wide vocabulary coverage; quantitative metrics are explored in Appendix A.3.
@@ 3.2 SmolDoc: LLMs with prompt mesh
SmolDoc follows a different and complementary approach. Whereas SmolSent consists of a small set of sentences that are selected from natural text, are complex, and cover many tokens; SmolDoc instead consists of documents that are generated and are simpler, but cover many topics. It should be noted that the token-coverage approach described above failed resoundingly for longer documents, as the power of the honeypots was magnified.
To generate SmolDoc, we used a collection of templates to create a few thousand diverse prompts with a wide range of topics, domains, words, tenses, grammatical cases, and registers (e.g. formal, informal). Appendix B.2 gives details and examples.
@@
Document-by-document evaluation as described above does not help one understand corpus diversity — for example, if an almost identical document appears twice, only one of them should be included. Therefore, we rank all candidates by how much new information they add to the corpus, by iteratively finding the document contributing the least new information and removing it, yielding a ranking of all documents. Our criterion for “new information” was the average 9-gram IDF score of a document, minus the fourth moment Bread score (Caswell et al., 2023), to down-weight internally repetitive documents.
@@
We wanted to translate text data for languages with more speakers. We break the languages into the five different groups/tiers seen in Appendix Table 6, and for each tier translate the top N documents as ranked by corpus diversity. The smallest tier has 66 documents, and the largest has 584.
3.3 Non-English-centric translations
For SmolDoc, we additionally collected data for four non-English-centric language pairs, namely from each of the East African languages of Amharic (am) and Swahili (sw) to each of regionally relevant languages Standard Arabic (ar) and Mandarin Chinese (zh). Including the reversed versions of these, this yields 8 total non-English-centric language pairs. Because of the difficulties of choosing/generating good source material in these languages, we used the human translations from English as the source text. However, due to the lack of appropriate evaluation sets, it is difficult to know the value-add of this data over datasets pivoted through English.
@@ 4 Data collection and verification
[…]
set | N langs | ex | tok | char | ex/LP | tok/LP | char/LP |
Gatitos | 176 | 693k | 784k | 4.6M | 3.9k | 4.5k | 26k |
SmolSent | 81 | 70k | 994k | 6.1M | 863 | 12k | 75k |
SmolDoc | 100 | 27k | 5.1M | 28M | 263 | 50k | 278k |
[…]
5 Finetuning
To get a general understanding of how helpful these data are, we fine-tune the Gemini 2.0 Flash model (Team et al., 2024a, b) on these data. It is important to note that the experimental set-up in this work is not optimized, and simply evaluates the “out of the box” improvement of simplistic finetuning with no bells and whistles. The ChrF scores could certainly be improved with e.g. a smaller set of languages, mixing in pre-training data, using larger base models, etc..
ChrF on LPs in… | SmolSent (80 LP) | SmolDocSp. (73 LP) | Overlap (38 LP) | |||
Model | 0-shot | 10-shot | 0-shot | 10-shot | 0-shot | 10-shot |
G. Translate | - | - | - | - | 43.2 | - |
NLLB-54b | - | - | - | - | 40.0 | - |
Claude 3.5 Son. | 37.5 | 39.7 | 38.3 | 40.9 | 41.0 | 42.8 |
GPT-4o | 29.9 | 34.1 | 31.8 | 36.3 | 35.4 | 38.5 |
Gemini 2.0 Pro | 38.9 | 38.9 | 39.9 | 40.3 | 42.6 | 42.2 |
[…]
+ SmolSent | 38.3 | 38.3 | 38.8 | 38.8 | 40.6 | 40.6 |
+ SmolDoc | 35.3 | 35.3 | 39.5 | 39.5 | 41.2 | 41.2 |
+ Both | 38.9 | 38.9 | 40.5 | 40.5 | 41.8 | 41.8 |
+ BothG | 39.4 | 39.3 | 41.0 | 40.9 | 42.1 | 42.2 |
[…]

### v2

@@ Abstract
We open-source Smol (Set of Maximal Overall Leverage),111
google/smol a suite of
training data
to unlock machine translation for low-resource languages. Smol has been translated into 124 (and growing) under-resourced languages (125 language pairs),222Experiments are mostly on a subset of 115 languages, before volunteer translations of additional languages finished. The paper title reflects this. including many for which there exist no previous public resources, for a total of 6.1M translated tokens. Smol comprises two sub-datasets, each carefully chosen for maximum impact given its size: SmolSent, a set of sentences chosen for broad unique token coverage, and SmolDoc, a document-level resource focusing on a broad topic coverage.
They join the already released Gatitos for a trifecta of paragraph, sentence, and token-level content.
We demonstrate that using Smol to prompt or fine-tune Large Language Models yields robust ChrF improvements. In addition to translation, we provide factuality ratings and rationales for all documents in SmolDoc, yielding the first factuality datasets for most of these languages.
@@
Isaac Caswell∗ Elizabeth Nielsen∗ Jiaming Luo Colin Cherry Geza Kovacs Hadar Shemtov Partha Talukdar Dinesh Tewari Baba Mamadi Diane Djibrila Diane Solo Farabado Cissé Koulako Moussa Doumbouya Edoardo Ferrante Alessandro Guasoni Christopher Homan Mamadou K. Keita Sudhamoy DebBarma Ali Kuzhuget David Anugraha Muhammad Ravi Shulthan Habibi Genta Indra Winata Anthony Munthali Sina Ahmadi Andrei Chemyshev Mingfei Lau Jonathan Eng Google {Research, Deepmind} NKo USA INC Stanford University Conseggio pe-o patrimònio linguistico ligure Universitas Indonesia University of Zurich Rochester Institute of Technology Capital One tyvan.ru Mari Institute of Language, Literature and History Led volunteer contributions for NKo, Ligurian, Zarma, Mooré, Kokborok, Russian, Indonesian, Hawrami+Gilaki+Sorani, Tumbuka, Meadow Mari, and Cantonese, respectively ∗{icaswell, eknielsen}@google.com
@@ 1 Introduction
There exist no professionally-translated data for most of the world’s 7000 or so languages, rendering tasks like Machine Translation near impossible. High-quality data is needed. However, it is not clear how best to use a limited budget for an expensive task like professional translation. As shown by the Gatitos dataset (Jones et al., 2023), word-level translations provide large benefits to translation quality for low-resource languages at the lowest cost. However, gains quickly saturate, as single tokens are not very expressive. Sentence-level data is better for a model once token-level data saturates, but it has much more inherent redundancy; and document-level data is even more effective…and more redundant.
In this work, we release the Smol dataset, which provides professionally translated sentence- and document-level data for 124 LRLs (125 language pairs). Smol contains two sub-datasets:
SmolSent: 863 English sentences covering 5.5k of the most common English tokens,333In this paper, ‘token’ refers to typographic units as an approximation to words, not subword tokens from a model’s vocabulary. professionally translated into 88 languages.
SmolDoc 584 English documents covering a wide range of topics, domains, and tokens, generated by an LLM and professionally translated into 106 languages.
We demonstrate the utility of these data for finetuning and prompting LLMs for translation, and provide factuality annotations for all documents.
@@ 2 Related work
There are not many training datasets with human-translated data for Low-Resource Languages (LRLs), where we operationally define LRL as any language beyond the first 100 supported by most traditional crawls and MT providers (enumerated in Appendix section A).
@@
Tatoeba (Tiedemann, 2020) is probably the most multilingual, but it is made of volunteer contributions and of unclear quality. The Gatitos dataset Jones et al. (2023) consists of a 4000-entry lexicon translated into 170 LRLs, but is only token-level. Most similar to the present work, NLLB-Seed is a high-quality, sentence-level training set of 6k sentences selected from English Wikipedia and professionally translated into 44 LRLs Team et al. (2022). There are also several professionally-translated evaluation sets, namely Flores-101 and Flores-200 Goyal et al. (2022); Team et al. (2022), and NTreX (Federmann et al., 2022a).
@@
While highly multilingual, professionally translated training data is rare, there is a growing number of bottom-up community data sources organized through research collectives like Masakhane ( et al., 2020), Turkish Interlingua (Mirzakhalov et al., 2021a, b), and GhanaNLP (Azunre et al., 2021a); and conferences and workshops like AfricaNLP, AmericasNLP (Mager et al., 2021) and ArabNLP. These datasets are usually generated by researchers fluent in the languages, and are therefore especially high quality. In addition to providing datasets, such efforts frequently also provide models and baselines, or even public interfaces, like the Khaya Translator Web App444https://ghananlp.org/project/translator-webapp/ by GhanaNLP for West African languages, and the lesan.ai555https://lesan.ai/translate translation website for Ethiopian languages.
@@
Participation is especially strong from the African continent, including corpora and models for pan-East-African languages (Babirye et al., 2022), languages from the Horn of Africa (Hadgu et al., 2022), Ethiopian languages (Teferra Abate et al., 2018; Gezmu et al., 2021), Ugandan languages (Akera et al., 2022), Emakhuwa (Ali et al., 2021), South-African languages (Eiselen and Puttkammer, 2014), Setswana and Sepedi (Marivate et al., 2020), Yorùbá (Adelani et al., 2021b, a), Oshiwambo (Nekoto et al., 2022), Igbo (Ezeani et al., 2020), Zulu (Mabuya et al., 2021), Twi (Azunre et al., 2021b), Gbe (Hacheme, 2021), Bambara (Tapo et al., 2021), and Fon (Emezue and Dossou, 2020). Outside of Africa, corpora have been created for languages of the Americas, including for four indigenous languages of Peru in Bustamante et al. (2020), the numerous results on the largely South- and Central American languages from the first AmericasNLP conference (Mager et al., 2021), and the Inuktitut language of Canada (Joanis et al., 2020). Datasets for lower-resourced languages of India have also sprung up, including the 13-language PMIndia (Haddow and Kirefu, 2020), and datasets focused on languages of the Northeast like Mizo (Thihlum et al., 2020), Khasi (Laskar et al., 2021) and Assamese (Laskar et al., 2020). Further West, PARME (Ahmadi et al., 2025) has provided some of the first human-translated content for Kurdish and Iranian languages. Finally, a variety of such datasets and models are available for public use on HuggingFace666https://huggingface.co/datasets?multilinguality=multilinguality:translation&task_categories=task_categories:translation or Zenodo.777https://zenodo.org/communities/africanlp/
3 Text Selection
Translation requires significant investment and can’t be easily re-done, so great care needs be put into carefully choosing sentences to translate. For both sub-datasets SmolDoc and SmolSent, selection or generation of source text is done in English. Selecting only English has clear biases, but also has advantages—most notably, for N languages, it requires N times less work to quality control. Future work should consider focusing on non-English sources.
@@ 3.1 SmolSent: Token Set Cover
Our basic motivation for creating SmolSent was to help models overcome vocabulary issues, which are common for the lowest-resource languages (Nielsen et al., 2025; Bapna et al., 2022). Therefore, we frame this as a set-cover problem, and pick the smallest set of sentences (from CommonCrawl888https://commoncrawl.org/ we use all available snapshots as of August 20, 2022) that covers the largest set of target tokens. The tokens we chose to cover (the target set) were the English side of Gatitos, as well as the most common 2,500 tokens from an English web crawl. Set cover is NP-hard, so we approximate it with a greedy algorithm that iteratively picks the sentence with the highest coverage percent, defined as the percentage of its tokens that are in the target set.
Preliminary work on Token Set-Cover
@@
To evaluate the token set-cover approach, we started by selecting data from existing web-scraped parallel data. We pretrain a multilingual Neural Machine Translation (NMT) model on parallel data from 294 language pairs from Madlad-400 Kudugunta et al. (2023), with nine languages held out to simulate LRLs. We fine-tune this model on sets of existing parallel data in each of the held-out languages, and evaluate on Flores-200. Details on the experimental set-up in Appendix B.1.
In addition to Greedy Token Set-Cover, we explore two methods that balance data diversity and data quality. First, we implement Ambati et al. (2011)’s ‘density-weighted diversity’ (DWD) metric, which is an -gram based metric for diversity and quality. Second, we implement an embedding-based version of DWD, which takes the weighted harmonic mean of perplexity under the Palm 2 model Anil et al. (2023) (proxy for quality), and embedding distance on mBERT sentence embeddings (proxy for diversity). We apply both methods to the English side of the parallel data only, to simulate the case where we don’t yet have LRL translations. As a baseline, we randomly select sentences.
@@
Researcher in the Loop (RITL)
Despite its success in the ablation, Greedy Token Set Cover had several problems when we scaled it to select from among all the English sentences of CommonCrawl. Firstly, it is maximized by honeypots, or nonsense strings dense in content words (Appendix Table B.1); and secondly, it biases towards short sentences, causing length distribution artifacts.
These problems are not easy to solve with heuristics—for example, if you disqualify lists with commas you’ll get ones with spaces, if you require sentences to have some function words or token-length diversity, you’ll get other sorts of garbled sentences, and so on. However, a dataset like Smol is small enough to manually inspect. Therefore we develop Researcher in the Loop Greedy Set-Cover (Algorithm 1), where the domain expert (the researcher) can look at and edit each individual sentence.999This work was conducted before the advent of LLMs. Today, this could be simplified using LLMs as autoraters. The result of this process is SmolSent, a set which uses 863 sentences to cover 5519 unique tokens. Qualitatively, SmolSent consists of complex sentences with wide vocabulary coverage; quantitative metrics are explored in Appendix B.3.
@@ 3.2 SmolDoc: LLMs with prompt mesh
SmolDoc follows a different and complementary approach. Whereas SmolSent consists of a small set of sentences that are selected from natural text, are complex, and cover many tokens; SmolDoc instead consists of documents that are generated and are simpler, but cover many topics. It should be noted that the token-coverage approach described above failed resoundingly for longer documents, as the prevalence of the honeypots was magnified.
To generate SmolDoc, we used a collection of templates to create a few thousand diverse prompts with a wide range of topics, domains, words, tenses, grammatical cases, and registers (e.g. formal, informal). Appendix C.2 gives details and examples.
@@
Document-by-document evaluation as described above does not help one understand corpus diversity—for example, if an almost identical document appears twice, only one of them should be included. Therefore, we rank all candidates by how much new information they add to the corpus, by iteratively finding the document contributing the least new information and removing it, thus ranking all documents. Our criterion for “new information” was the average character 9-gram Inverse Document Frequency (IDF) score of a document—in other words, how rare its substrings were across all of the documents in the pool so far. To down-weight internally repetitive documents, we substracted the fourth moment Bread score (Caswell et al., 2023).
@@
We wanted to translate more data for languages with more speakers. We break the languages into the five different groups, each with a larger subset of the generated documents. Each tier contains translations of the top N documents as ranked by corpus diversity. These can be seen in Appendix Table C.1.
Non-English-centric translations
For SmolDoc, we additionally collected data for four non-English-centric language pairs, from each of the East African languages of Amharic (am) and Swahili (sw) to each of regionally relevant languages Standard Arabic (ar) and Mandarin Chinese (zh). Including the reversed versions of these, this yields 8 total language pairs. Because of the difficulty of generating good source material in these languages, we used the existing SmolDoc translations to Swahili/Amharic as the source text. However, due to the lack of appropriate evaluation sets, it is hard to know the value-add of this data over datasets pivoted through English.
@@ 4 Data Collection and Verification
Several languages are contributed by volunteers; they are listed as co-authors.101010Community contributions of translations or corrections are welcome; please reach out to the authors or join the TUSL Discord. For the other languages, the translation provider we contracted has worked with us for many years, and has a pre-existing relationship with professional translators for all languages in the Smol datasets. The translators are paid a fair wage, and their identities are contractually kept anonymous to us. We checked the delivery for duplicate translations, anomalous source/target length ratios, and similarity with Google Translate outputs. Very few languages were flagged this way. Following this, we ran FunLangID (Caswell, 2024) on all segments and found no issues. Manual inspection turned up several issues with nonunicode fonts (e.g. ô for =) for West African languages, and nonstandard orthography for Santali; these issues were then fixed. The choice of script, orthography and translation variety was challenging for many communities, including Kurdish, Zaza-Gorani and Gilaki languages, all of which have more than one orthography and lack a standard variety.
The largest missing check is for fluency, which is hard to measure without trusted native speakers outside of the translation agency, or trusted LLMs; neither of which exist for all Smol languages.
| Total Dataset | Per Language Pair (LP) | ||||||
| Set | # languages | Examples | Tokens | Characters | Examples | Tokens | Characters |
| Gatitos | 178 | 714k | 807k | 4.8M | 4.0k | 4.5k | 27k |
| SmolSent | 88 | 76k | 1.1M | 6.6M | 863 | 12k | 75k |
| SmolDoc | 106 | 27k | 5.1M | 29M | 256 | 48k | 270k |
| Smol{Sent+Doc} | 125 | 103k | 6.2M | 35M | 824 | 50k | 282k |
| LP subset | Smol-Sent (80 LP) | Smol-Doc (73 LP) | Intersect (38 LP) | Hard (32 LP) | ||||
| Model | 0-shot | 10-shot | 0-shot | 10-shot | 0-shot | 10-shot | 0-shot | 10-shot |
| G. Translate | - | - | - | - | 43.2 | - | - | - |
| NLLB-54b | - | - | - | - | 40.0 | - | - | - |
| Claude 3.5 Son. | 37.5 | 39.7 | 38.3 | 40.9 | 41.0 | 42.8 | 30.0 | 33.5 |
| GPT-4o | 29.9 | 34.1 | 31.8 | 36.3 | 35.4 | 38.5 | 15.9 | 23.7 |
| Gemini 2.0 Pro | 38.9 | 38.9 | 39.9 | 40.3 | 42.6 | 42.2 | 31.4 | 31.7 |
| Gemini-2.0 Flash | 35.6 | 38.4 | 36.9 | 39.7 | 40.2 | 41.4 | 26.3 | 30.4 |
| + SmolSent | 38.3 | 38.3 | 38.8 | 38.8 | 40.6 | 40.6 | 32.5 | 32.6 |
| + SmolDoc | 35.3 | 35.4 | 39.5 | 39.5 | 41.2 | 41.2 | 31.8 | 31.8 |
| + Both | 38.9 | 38.9 | 40.5 | 40.5 | 41.8 | 41.8 | 33.4 | 33.4 |
| + Both+G | 39.4 | 39.3 | 41.0 | 40.9 | 42.1 | 42.2 | 33.9 | 33.9 |
| +3.8 | +0.9 | +4.1 | +1.2 | +1.9 | +0.8 | +7.6 | +3.5 | |
5 Finetuning and In-Context Learning
We use fine-tuning and ICL as tools to demonstrate the value of the Smol dataset. As this is a data paper, these experiments are motivated by the maxim “what could any researcher simply train with public APIs?” More involved techniques, e.g. Reinforcement-Learning-based approaches, will likely lead to stronger results.